# Overflow Prevention Enhances Long-Context Recurrent LLMs

### v1

@@ 5.1 Long-Context evaluations
LongBench.  LongBench (Bai et al., 2024) is an extensive real-world long-context benchmark that combines 16 different tasks across 6 categories: Single-Document QA, Multi-Document QA, Summarization, Few-Shot Learning, Synthetic Tasks, and Code Completion. Here, most models do not suffer from limited length generalization, as they were trained on contexts of same or longer length. Results for leading recurrent LLMs with and without OPRM inference are presented in Figure 1 (right). Full results can be found in Tables 9101112 in the Appendix. By allocating a different state per chunk, OPRM is able to mitigate memory overflows. This change proves highly effective, as OPRM improves performance in almost all categories. The total improvement when applying OPRM inference is 14% for Falcon3-Mamba-Inst-7B, 28% for Falcon-Mamba-Inst-7B, 50% for Recurrent-Gemma-IT-9B, and 51% for RWKV6-Finch-7B. Additionally, we observe that as context length increases, the advantage of OPRM becomes more evident (4-8k, 8k+). This is very reasonable, as the amount of information in the context correlates with its length.
@@ 5.2 Context Extension
Here, we focus on models that were trained on short sequences, and evaluate them on long sequences. As OPRM naturally performs context extension, we compare it to dedicated methods, and show that the latter are less effective, as they do not prevent overflows. An additional Needle-In-A-Haystack extension experiment is provided in Appendix 8.
@@ 6 Limitations
While our overflow-prevention algorithm yields significant gains, it can still be improved. For example, integrating cross-chunk processing would allow for better utilization of global in-context dependencies. Additionally, OPRM is training-free, therefore it relies on the trained model’s abilities. E.g., the IDK filter could be better adapted to Falcon3-Mamba-Inst-7B via additional fine-tuning.
@@ A.5 Document Retrieval
We use the same setup as in Ben-Kish et al. (2025). We train each model with data from SQuAD v2 (Rajpurkar et al., 2018), which provides examples in the form of (Query, Document, Answer). The training samples have the following form: $N_{docs}\times$$<$$Document$$>$; $<$$Answer$$>$, where $<$$Document$$>$ can be either the golden document (which holds the answer to the query) or one of $N_{docs}-1$ randomly sampled documents. $<$$Answer$$>$ holds the id of the golden document. In this setting $N_{docs}=11$, the order of the documents is random, and the query and respective document id are appended to the beginning of each document. During evaluation the same setting is usedm but the value of $N_{docs}$ is varied between 11 and 240. (between 2,200 tokens to 50,000 tokens). We train the 1.4b models for 400 steps, use a learning rate of 2e-5, gradient clipping of 1, batch size of 64 (used batch accumulation), and AdamW optimizer with weight decay of 0.1. For DeciMamba we use decimation_layer = 11, $L_{base}$=2000 during training and $L_{base}$=5000 during evaluation. We apply OPRM inference on top of the baseline model, and use chunk size $L=6000$. We train the 130m models for two epochs (1500 steps in each), use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation), and AdamW optimizer with weight decay of 0.1. For DeciMamba we use decimation_layer = 12, $L_{base}$=2000 during training and $L_{base}$=4000 during evaluation. We apply OPRM inference on top of the baseline model, and use chunk size $L=4000$. All results are the average performance over 3 different training seeds, and each data point is evaluated over 100 samples.
@@ Appendix B Additional Experiments
B.1 Efficiency of OPRM
In Table 8 we report the inference time and memory usage of a Falcon3-Mamba-Inst-7B model with and without OPRM inference (pre-fill + decoding of 10 tokens) on a Nvidia RTX A6000 GPU. In this test we use a chunk size $L=2000$ tokens. We find that as context lengths increase, our method outperforms the baseline model in terms of speed. This is not surprising, as OPRM inference enables additional parallelization in the sequence dimension (chunking), and reduces complexity, as detailed in Appendix D. Furthermore, we find that OPRM ’s memory usage is highly competitive w.r.t the baseline model, despite using much more states. For example, when the context length is 128K tokens, we use 64 chunks, which requires storing 63 more states in the GPU’s memory. The relatively small increase in memory usage is due to the fact that the memory occupied by a single state is about three orders of magnitude smaller than the size occupied by the model’s weights. This allows us to significantly increase the number of states (chunks) with only a slight increase in total memory usage.
Model Context Length
2K 4k 8K 16K 32K 64K 128K
Time [s] Falcon3-Mamba-Inst-7B + OPRM 2.2 2.5 3.2 4.7 7.8 14.0 26.9
Falcon3-Mamba-Inst-7B 2.0 2.5 3.5 5.7 10.2 18.9 36.2
Falcon3-Mamba-Inst-7B 14.9 15.3 15.9 17.2 19.9 25.2 35.7
Table 8: Efficieny of OPRM. Inference time (seconds) and peak memory usage (GigaBytes) benchmarked on a Nvidia RTX A6000 GPU. For OPRM inference we use $L=2000$.
@@ B.1 Efficiency of OPRM
Needle in a Haystack.  Following the setting in Ben-Kish et al. (2025), a Mamba-130m model needs to retrieve a random 5-digit code (needle) hidden at a random location within a concatenation of articles (haystack) sampled from WikiText (Merity et al., 2016). While the base model was trained on context lengths of 2K tokens, during inference we increase the sequence lengths exponentially from 1K to 512K and record the model’s performance for a variety of needle depths within the context. Results can be found in Figure 4. We find that OPRM extends the context to sequences that are ×$\times 256$ longer than those seen during training. This is in contrast to the dedicated method, which is able to extend the context by ×$\times 64$. The baseline model (Mamba) is able to extend the context by ×$\times 8$ (not shown).
@@ Appendix D Advantages of OPRM
Assuming that the lengthy part of the prompt $X$ is the context $C$ $(|C|\gg|P|,|S|)$, OPRM offers two key efficiency advantages over other LLMs: (i) OPRM processes each chunk independently and thus reduces complexity. As most recurrent LLMs employ FFT-based methods or work-efficient parallel scan algorithms during prefill, both leading to an $O(Lb\log Lb)$ aggregated time-complexity for processing a sequence of length $Lb$, OPRM improves efficiency by reducing the prefill time complexity to $O(bL\log L)$. Additionally, (ii) in many cases, the context C and prefix P are known in advance, while the query (suffix) is provided in real time. OPRM can leverage this by pre-computing the recurrent states for each chunk of the context $[P,C_{i}]$. At inference time, instead of recomputing from scratch, the model initializes with these precomputed states, allowing efficient real-time query processing and significantly reducing computational overhead, resulting in a prefill complexity of $O(b|S|)$ via sequential prefill, which isn’t dependent on the context length $|C|$. Note that such an optimization cannot be applied to Transformers, as they are not recurrent.
@@
Although our method can be applied to any long-text scenario, as the Prefix-Context-Suffix structure is a general framework, it is also well-suited for query-based RAG (Lewis et al., 2020). This setting focuses on dynamically retrieving relevant context based on the query before feeding it into the LLM. Our approach is particularly efficient in this setting, as it processes the query with the selected chunk $C_{j}$ in a single pass performed when computing the state $S_{j}$. This makes it a simple yet effective drop-in component for various tasks and real-world applications. While our approach shares similarities with RAG methods, we emphasize that our primary focus is not on direct comparisons with RAG benchmarks. Instead, our goal is to enhance recurrent LLMs’ long-context capabilities.

### v2

@@ 5.1 Long-Context evaluations
LongBench.  LongBench (Bai et al., 2024) is an extensive real-world long-context benchmark that combines 16 different tasks across 6 categories: Single-Document QA, Multi-Document QA, Summarization, Few-Shot Learning, Synthetic Tasks, and Code Completion. Here, most models do not suffer from limited length generalization, as they were trained on contexts of same or longer length. Results for leading recurrent LLMs with and without OPRM inference are presented in Figure 1 (right). Full results can be found in Tables 13141516 in the Appendix. By allocating a different state per chunk, OPRM is able to mitigate memory overflows. This change proves highly effective, as OPRM improves performance in almost all categories. The total improvement when applying OPRM inference is 14% for Falcon3-Mamba-Inst-7B, 28% for Falcon-Mamba-Inst-7B, 50% for Recurrent-Gemma-IT-9B, and 51% for RWKV6-Finch-7B. Additionally, we observe that as context length increases, the advantage of OPRM becomes more evident (4-8k, 8k+). This is very reasonable, as the amount of information in the context correlates with its length. Lastly, as shown in Table 12 (Appendix), we find that OPRM significantly improves multi-hop reasoning performance, which is more than doubled on certain benchmarks across all evaluated models. These results suggest that memory overflows severely degrade multi-hop reasoning capabilities, and that OPRM provides an efficient and effective solution.
@@ 5.2 Context Extension
Here, we focus on models that were trained on short sequences, and evaluate them on long sequences. As OPRM naturally performs context extension, we compare it to dedicated methods, and show that the latter are less effective, as they do not prevent overflows. An additional Needle-In-A-Haystack extension experiment is provided in Appendix B.3.
@@ 5.3 Ablations
5.4 Efficiency of OPRM
In Table 8 we report the inference time and memory usage of a Falcon3-Mamba-Inst-7B model with and without OPRM inference (pre-fill + decoding of 10 tokens) on a Nvidia RTX A6000 GPU. In this test we use a chunk size of $L=2000$ tokens. We find that as context length grows, OPRM becomes increasingly faster than the baseline. This is not surprising, as OPRM allows additional parallelization in the sequence dimension (chunking), and reduces complexity, as detailed in Appendix D. Furthermore, we find that OPRM ’s memory usage is highly competitive w.r.t the baseline model, despite using much more states. For example, when the context length is 128K tokens, we use 64 chunks, which requires storing 63 more states in the GPU’s memory. The relatively small increase in memory usage is due to the fact that the memory occupied by a single state is about three orders of magnitude smaller than the size occupied by the model’s weights. This allows us to significantly increase the number of states (chunks) with only a slight increase in total memory usage.
Model Context Length
2K 4k 8K 16K 32K 64K 128K
Time [s] Falcon3-Mamba-Inst-7B + OPRM 2.2 2.5 3.2 4.7 7.8 14.0 26.9
Falcon3-Mamba-Inst-7B 2.0 2.5 3.5 5.7 10.2 18.9 36.2
Falcon3-Mamba-Inst-7B 14.9 15.3 15.9 17.2 19.9 25.2 35.7
Table 8: Efficieny of OPRM. Inference time (seconds) and peak memory usage (GigaBytes) benchmarked on a Nvidia RTX A6000 GPU. For OPRM inference we use $L=2000$. We find that OPRM outperforms vanilla inference in speed, while adding a surprisingly small memory overhead.
@@ 6 Limitations
While our overflow-prevention algorithm yields significant gains, it can still be improved. For example, non-trivial overflow-aware cross-chunk aggregation methods would allow for better utilization of global in-context dependencies. Additionally, OPRM is training-free, therefore it relies on the trained model’s abilities. E.g., the IDK filter could be better adapted to Falcon3-Mamba-Inst-7B via additional fine-tuning.
@@ A.5 Document Retrieval
We use the same setup as in Ben-Kish et al. (2025). We train each model with data from SQuAD v2 (Rajpurkar et al., 2018), which provides examples in the form of (Query, Document, Answer). The training samples have the following form: $N_{docs}\times$<<DocumentDocument>>; <<AnswerAnswer>>, where <<DocumentDocument>> can be either the golden document (which holds the answer to the query) or one of $N_{docs}-1$ randomly sampled documents. <<AnswerAnswer>> holds the id of the golden document. In this setting $N_{docs}=11$, the order of the documents is random, and the query and respective document id are appended to the beginning of each document. During evaluation the same setting is usedm but the value of $N_{docs}$ is varied between 11 and 240. (between 2,200 tokens to 50,000 tokens). We train the 1.4b models for 400 steps, use a learning rate of 2e-5, gradient clipping of 1, batch size of 64 (used batch accumulation), and AdamW optimizer with weight decay of 0.1. For DeciMamba we use decimation_layer = 11, $L_{base}$=2000 during training and $L_{base}$=5000 during evaluation. We apply OPRM inference on top of the baseline model, and use chunk size $L=6000$. We train the 130m models for two epochs (1500 steps in each), use a learning rate of 1e-4, gradient clipping of 1, batch size of 64 (used batch accumulation), and AdamW optimizer with weight decay of 0.1. For DeciMamba we use decimation_layer = 12, $L_{base}$=2000 during training and $L_{base}$=4000 during evaluation. We apply OPRM inference on top of the baseline model, and use chunk size $L=4000$. All results are the average performance over 3 different training seeds, and each data point is evaluated over 100 samples.
@@ Appendix B Additional Experiments
B.1 InfiniteBench
InfiniteBench (Zhang et al., 2024) comprises both synthetic and real-world tasks across diverse domains, designed to evaluate a model’s ability to understand long-range dependencies within extended contexts exceeding 100k tokens.
Table 9 shows a comparison between Falcon3-Mamba-Inst-7B with and without OPRM, and additional Attention-based models. Consistent with our earlier findings, OPRM significantly outperforms the baseline, and even beats the equivalently-sized Attention-based model, Mistral-YaRN-7B . Moreover, with OPRM, Falcon3-Mamba is competitive with Attention-based models that have more than x10 parameters such as Claude2 and Kimi-Chat. Lastly, we note that Code.Run is not solved by any model (except GPT4, to some extent). This synthetic task - which requires computing a complex composition of many functions - remains unsolved even by leading proprietary models. The same is true for Math.Calc.
Model Kimi-Chat Claude2 GPT4 Mistral-YaRN Falcon3-Mamba Falcon3-Mamba + OPRM
Model Type Transformer Transformer Transformer Transformer Recurrent Recurrent
# Params >>70B >>70B >>70B 7B 7B 7B
Ret.PassKey 98.1 97.8 100.0 92.7 0.0 99.8
Ret.Number 95.4 98.1 100.0 56.6 0.0 100.0
Ret.KV 53.6 65.4 89.0 0.0 0.0 31.3
En.Sum 17.9 14.5 14.7 9.1 20.1 22.08
En.QA 16.5 12.0 22.2 9.6 11.0 23.2
En.MC 72.5 62.9 67.3 28.0 45.4 59.4
En.Dia 11.5 46.5 8.5 7.5 4.0 8.0
Code.Dbg 18.0 2.3 39.6 0.8 27.1 24.56
Code.Run 2.0 2.5 23.3 1.3 0.0 0.0
Math.Calc 0.0 0.0 0.0 0.0 0.0 0.0
Math.Find 12.6 32.3 60.0 17.1 26.86 33.14
Avg 36.2 39.5 47.7 20.2 12.2 36.5
Table 9: InfiniteBench. Results for Falcon3-Mamba-Inst-7B, with and without OPRM inference, along with Attention-based LLMs. Large proprietary models with a parameter count larger by more than 10 times are added in gray. We find that with OPRM, leading recurrent LLMs, such as Falcon3-Mamba-Inst-7B, can outperform similar sized Transformer-based models, and are comparable to larger proprietary models.
B.2 Comparison to RAG
RAG methods augment LLMs with information retrieved from a large input source, such as a large document database (Lewis et al., 2020). Since RAG methods can be easily applied to long-context tasks, we compare them to OPRM. As OPRM is training-free, we specifically compare it to two zero-shot RAG methods: DRAGON, which uses a dense retriever designed to generalize in zero-shot settings (Lin et al., 2023), and PRP, a retriever-free approach that ranks chunks using additional forward passes (Qin et al., 2024a). We evaluate these methods across all four Document QA tasks in LongBench_e (HotpotQA, 2WikiMQA, MultiFieldQA, and Qasper), and report the average performance across all benchmarks. For each RAG method, we tested multiple combinations of num_chunks and chunk_size, ensuring a fair comparison by constraining num_chunks x chunk_size = OPRM_chunk_size. As shown in the table below, OPRM outperforms both methods:
Method 0-4K 4K-8K 8K+ Avg
Falcon-Mamba-Inst-7B 34.21 26.94 19.11 26.75
+ Dragon 35.85 30.05 32.74 32.88
+ PRP 36.27 32.52 34.61 34.47
+ OPRM 37.41 34.49 36.25 36.05
Table 10: Comparison to RAG Methods. We compare the above methods across all four Document QA tasks in LongBench_e and report the average score. 0-4K, 4-8K, and 8k+ are the LongBench_e length groups, and Avg is the overall average score. While RAG methods augment the baseline model, we find that OPRM outperforms them, especially as context lengths increase.
Our results are consistent with prior findings: RAG-based methods generally do not outperform long-context LLMs on long-context tasks (Li et al., 2024; Bai et al., 2024; Xu et al., 2024). To the best of our knowledge, there is no definitive explanation - yet, one hypothesis suggests that longer contiguous contexts help the model uncover multi-hop relations, as some reasoning steps depend on access to previously seen information (Xu et al., 2024). In such cases, given the same length budget, selecting longer segments may be more effective than using multiple smaller chunks: when a chunk is small, its relevance is more likely to be biased toward surface-level similarity with the query. In contrast, longer segments include additional context that can improve the relevance estimation.
Moreover, from an efficiency perspective - the recurrent LLMs used in our work share the same computational complexity as RAG-based solutions. This stands in contrast to the Transformer-based long-context LLMs used in (Li et al., 2024; Bai et al., 2024; Xu et al., 2024), which incur a substantial computational overhead. Lastly, it is also worth noting that PRP’s sequential ranking process makes it significantly slower than both OPRM and DRAGON: PRP requires several minutes to process each sample, whereas OPRM and DRAGON take only a few seconds.
To conclude, these findings further underscore the potential of long-context recurrent LLMs when combined with overflow prevention techniques such as OPRM.
B.3 Needle in a Haystack
@@
Following the setting in Ben-Kish et al. (2025), a Mamba-130m model needs to retrieve a random 5-digit code (needle) hidden at a random location within a concatenation of articles (haystack) sampled from WikiText (Merity et al., 2016). While the base model was trained on context lengths of 2K tokens, during inference we increase the sequence lengths exponentially from 1K to 512K and record the model’s performance for a variety of needle depths within the context. Results can be found in Figure 4. We find that OPRM extends the context to sequences that are $\times 256$ longer than those seen during training. This is in contrast to the dedicated method, which is able to extend the context by $\times 64$. The baseline model (Mamba) is able to extend the context by $\times 8$ (not shown).
@@
B.4 Multi-Chunk Ablation
To further motivate our single-chunk approach, we compare OPRM to a multi-chunk strategy. We use the OPRM pipeline to first identify the top-k relevant chunks and then evaluate them together with an additional forward pass. We refer to this method as CC (Combined Chunks). We note that all methods use the same token budget - CC uses smaller chunks with a total length equal to OPRM’s single chunk. We evaluated CC’s performance across all 4 Document QA tasks in LongBench_e and report the average scores in Table 11.
Method 0-4K 4K-8K 8K+ Avg
Falcon-Mamba-Inst-7B 34.21 26.94 19.11 26.75
+ CC 40.27 35.56 35.98 37.27
+ OPRM 37.41 34.49 36.25 36.05
Table 11: Multi-Chunk Ablation. We compare all methods on the 4 Document QA datasets in LongBench_e (HotpotQA, 2WikiMQA, MultiFieldQA, and Qasper) and report the average score for each length group. Falcon-Mamba-Inst-7B is the baseline (vanilla inference). ‘+’ indicates the inference algorithm. The total length of the $C$ chunks is constrained to equal one OPRM chunk. We experimented with multiple combinations of chunk length and number, and used the best parameters for testing.
We see that while CC is beneficial for short contexts (0-4k), it offers no clear advantage over OPRM in longer contexts (4-8k and 8k+). Moreover, it adds computational and algorithmic overheads. One possible explanation is that longer contiguous contexts help the model uncover multi-hop relations (Xu et al., 2024), as some reasoning steps depend on access to previously seen information. In such cases, given the same length budget, selecting longer segments may be better than using multiple smaller chunks: when a chunk is small, its relevance is more likely to be biased toward surface-level similarity with the query. In contrast, longer segments include additional context that can improve the relevance estimation.
We conclude that cross-chunk information fusion warrants further investigation. For example, a system-oriented approach, such as an overflow-aware variant of Bahdanau Attention (Bahdanau et al., 2016) or hierarchical state processing, may hold significant potential.
B.5 Sensitivity of Recurrent Memory Capacity to Input Length
We repeat the zero-shot AR experiment (Section 3) with varying sequence lengths and present the results in Figure 5. Our findings indicate that memory capacity is not particularly sensitive to the overall context length. Similar to the original 1,200-token setting, performance starts at around 75% accuracy for 10 key-value pairs (facts), and gradually declines as the number of facts increases, eventually converging toward 0% accuracy. This further demonstrates that recurrent memory overflows represent a fundamentally distinct limitation of recurrent LLMs - one that should be addressed independently of other known limitations, such as length generalization. Lastly, since each fact is 6 tokens long, for a sequence length of L=300 we can only test a maximum of 50 facts and for L=600 a maximum of 100 facts.
Figure 5: Sensitivity of recurrent memory capacity to input length. We repeat the zero-shot AR experiment (Section 3) with varying input sequence lengths $L\in\{300,600,1200,2400,4800\}$. We find that the model is not sensitive to the input sequence length, but rather to the amount of information within the sequence.
B.6 Positional Sensitivity to Recurrent Memory Overflows
We analyze the position of successfully retrieved key-value (KV) pairs from the zero-shot AR experiment (Section 3). The results are displayed in Figure 6. Here, each plot shows a normalized histogram of the locations of successfully retrieved KV pairs (facts). In all measurements we use 10 equally spaced bins to aggregate the samples, and average over 5 different seeds. When the number of facts is small, the distribution appears relatively uniform, which corresponds with the high AR accuracy. Interestingly, as the number of facts increases, we observe a U-shaped recall pattern, resembling the lost-in-the-middle phenomenon reported in (Liu et al., 2024a). We conclude that while recall performance degrades across all positions, the middle portions of the context suffer the most. Since this phenomenon is also observed in Transformer-based LLMs, we suspect that the positional sensitivity may be influenced by properties of the data itself rather than the model’s architecture, a finding which warrants further investigation in future work.
Figure 6: Positional sensitivity to recurrent memory overflows. We analyze the position of successfully retrieved key-value (KV) pairs from the zero-shot AR experiment (Section 3). Each plot shows a normalized histogram of the locations of successfully retrieved KV pairs (facts). We find that while recall performance degrades across all positions, the middle portions of the context suffer the most, resembling the ”Lost-In-The-Middle” phenomenon observed in Transformer-based LLMs.
B.7 Comparison to Agentic Long-Context Frameworks
One way to process long contexts is by applying an agentic pipeline that first splits the context into several chunks for LLMs to process and reason over, and then aggregates their intermediate answers using additional LLM calls to produce the final output. These methods contrast with inference methods like OPRM, which process the whole context in a single forward pass. We compare OPRM to LLMxMapReduce, an agentic framework for long context processing (Zhou et al., 2024). We apply LLMxMapReduce to Falcon3-Mamba-Inst-7B, and find that while it produces reasonable answers for some queries, it often fails at following the full set of instructions at different stages of the pipeline. Notably, in their paper, LLMxMapReduce use Transformer models with a minimum size of 70 billion parameters, whereas Falcon3-Mamba-Inst-7B, at 7 billion parameters, is currently the largest available recurrent LLM. We believe this performance gap is largely due to the significant difference in model scale, and conclude that it will be beneficial to evaluate LLMxMapReduce on larger-scale recurrent LLMs, once they become available.
Most importantly, we report that OPRM is over an order of magnitude faster than LLMxMapReduce: while LLMxMapReduce requires 225.2 seconds per query, OPRM completes a query in 8.6 seconds - a total speedup of x26 times.
@@ Appendix D Advantages of OPRM
Assuming that the lengthy part of the prompt $X$ is the context $C$ $(|C|\gg|P|,|S|)$, OPRM offers two key efficiency advantages over other LLMs: (i) OPRM processes each chunk independently and thus reduces complexity. As most recurrent LLMs employ FFT-based methods or work-efficient parallel scan algorithms during prefill, both leading to an $O(Lb\log Lb)$ aggregated time-complexity for processing a sequence of length $Lb$, OPRM improves efficiency by reducing the prefill time complexity to $O(bL\log L)$. Additionally, (ii) in many cases, the context C and prefix P are known in advance, while the query (suffix) is provided in real time. OPRM can leverage this by pre-computing the recurrent states for each chunk of the context $P,C_{i}]$. At inference time, instead of recomputing from scratch, the model initializes with these precomputed states, allowing efficient real-time query processing and significantly reducing computational overhead, resulting in a prefill complexity of $O(b|S|)$ via sequential prefill, which isn’t dependent on the context length $|C|$. Note that such an optimization will not be as effective for Transformers, as they do not compress the previously processed inputs.
@@
Although our method can be applied to any long-text scenario, as the Prefix-Context-Suffix structure is a general framework, it is also well-suited for query-based RAG (Lewis et al., 2020). This setting focuses on dynamically retrieving relevant context based on the query before feeding it into the LLM. Our approach is particularly efficient in this setting, as it processes the query with the selected chunk $C_{j}$ in a single pass performed when computing the state $S_{j}$. This makes it a simple yet effective drop-in component for various tasks and real-world applications. While our approach shares similarities with RAG methods, we emphasize that our primary focus is not on direct comparisons with RAG benchmarks. Instead, our goal is to enhance recurrent LLMs’ long-context capabilities. Furthermore, we find that our method outperforms several RAG methods, as shown and discussed in Section B.2.
Benchmark Method 0-4K 4K-8K 8K+ LB
HotPotQa (2 hops) Baseline 27.97 21.57 17.21 22.17
+ OPRM 38.68 34.37 36.07 35.09
Improvement 38.3% 59.3% 109.7% 58.3%
MuSiQue ($\leq 4$ hops) Baseline N/A N/A N/A 8.37
+ OPRM N/A N/A N/A 18.4
Improvement N/A N/A N/A 119.8%
2WikiMQA ($\leq 5$ hops) Baseline 25.26 25.33 16.61 21.39
+ OPRM 30.37 28.88 27.01 25.08
Improvement 20.2% 14.0% 62.7% 17.2%
Table 12: MultiHop Reasoning. We test SoTA recurrent LLMs with and without OPRM on all multi-hop benchmarks in LongBench, and report the average score over all models: Falcon3-Mamba-Inst-7B, Falcon-Mamba-Inst-7B, RecurrentGemma-IT-9B, and RWKV6-Finch-7B. Improvement shows the relative improvement in %. 0-4K, 4-8K, and 8k+ are the LongBench_e length groups, LB is the LongBench score. OPRM significantly improves multi-hop reasoning capabilities, sometimes by even more than 100%. The data is taken from Tables 13141516 in the Appendix.